OpenStudy (anonymous):

perform the indicated operation simplify if possible 8/3x-12 + x/x^2-16

OpenStudy (anonymous):

\[\frac{ 8 }{ 3x-12 } + \frac{ x }{ x^2-16 }\]

OpenStudy (dayakar):

\[\frac{ 8 }{ 3(x-4) }+\frac{ x }{ (x+4)(x-4) }\]

OpenStudy (dayakar):

can u do remaining part

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@dayakar i know the top would be 8x and the bottom would be 3(x+4)(x-4)

OpenStudy (dayakar):

\[\frac{ 8(x+4)+3x }{ 3(x+4)(x-4) }\] \[=\frac{ 8x+32+3x}{ 3(x+4)(x-4) }\] \[=\frac{ 11x+32 }{ 3(x+4)(x-4) }\]

OpenStudy (dayakar):

is it clear

OpenStudy (anonymous):

yea